What is the solution to (25x^2)-(16y^2)=400 ?
Answer by dabanfield(803)

(25x^2)-(16y^2)=400 ?
Divide by sides of the equation by 400:
25x^2/400 -16y^2/400 = 1
or
x^2/(400/25) - y^2/(400/16) = 1
This is in the standard form of a hyperbola [x^2/a^2 - y^2/b^2 = 1] which intersects the x-axis at (20/5,0) and (-20/5,0).
There are two asymtotes: y = (400/16)/(400/25)*x and y = -(400/16)/(400/25)*x